Split 6 liters of sugar water among 8 jars. Puts the same amount in each jar.how many milliliters of sugar water are in each jar
Answer by Alan3354(69443)

Split 6 liters of sugar water among 8 jars. Puts the same amount in each jar.how many milliliters of sugar water are in each jar
--------
6/8 = 3/4 liter per jar
= 750 ml per jar
============
6 liters = 6000 ml
6000/8 = 750 ml per jar